# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 6.2 Existence theorem
After making some minor modifications to McCann’s proof in [46, Section 6], we proceed by translating $\rho$ such that it enjoys a plane of symmetry $z=0$, and then use the results above to show $\rho$ is supported away from the boundary of $\Omega_{1-m}\cup\Omega_{m}$, then we finalize the proof of existence.
@@
Take $m$ small enough so that minimizers exist. First, we claim that any minimizer $\rho(m)=\rho_{m}+\rho_{1-m}$ for $E_{J}(\rho)$ on $W_{m}$ may be translated so that both $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=0$. In fact, since the $\Omega_{m}$ and $\Omega_{1-m}$ are convex and symmetric about $z=0$, it is enough to know that $\rho(m)$ enjoys a plane of symmetry $z=c$. This follows from a strong rearrangement inequality in Lieb [41, Lemma 3] and Fubini’s Theorem: by the definition of $U(\rho)$ and $I(\rho)$, they will not be changed after the symmetric decreasing rearrangement of $\rho$ along lines parallel to the z-axis; however, since $\left(r^{2}+z^{2}\right)^{-\frac{1}{2}}$ is strictly decreasing as a function of $|z|$, the rearrangement increases $G(\rho,\rho)$ unless $\rho$ is already symmetric decreasing about a plane $z=c$. Note $\Omega_{1-m}$ and $\Omega_{m}$ are convex, hence $\rho(m)$’s rearrangement is in $W_{m}$. Since $\rho(m)$ minimizes $E_{J}(\rho)$, then $G(\rho,\rho)$ cannot be increased after rearrangement, which means $\rho(m)$ enjoys a plane of symmetry $z=c$.
@@
Let $\delta$ as in Lemma 6.3 and $\rho(m)$ be the constrained energy minimizer in Theorem 4.14, then Lemma 6.3 shows that spt $\rho({m})$ is compact in the interior of $\Omega_{m}\cup\Omega_{1-m}$, therefore separated from the boundary by a positive distance $\zeta$. Lemma 2.9 (ii) shows that if $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho(m),\kappa)<\eta$, then $\kappa$ in $W_{m}$. Thus $E_{J}(\rho(m))\leq E_{J}(\kappa)$, and $\rho(m)$ is a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$ (Existence of $W^{\infty}$ local energy minimizer). By definition of $W_{m}$, we know all the functions in $W_{m}$ have compact support. Thus after translation we can set $\rho(m)\in{R}_{0}\left(\mathbb{R}^{3}\right)$. Since $\rho(m)$ is a $W^{\infty}$ local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$, in particular on ${R}_{0}\left(\mathbb{R}^{3}\right)$, Theorem 3.1 provides a local minimizer $(\rho(m),v)$ of $E(\rho,v)$ subject to the constraint on $J_{z}=J$, with $v(x)=$ $\frac{J}{I(\rho(m))}\hat{e}_{z}\times x$. [46, Section 2] and [15, Remark 2.31] show that $(\rho(m),v)$ minimizes $E(\rho,v)$ subject to the constraint on the vector angular momentum as well (Part (i)). Part (ii) comes from Lemma 6.3 (Part (ii)). By Theorem 3.1 we know $\rho(m)$ is continuous and satisfies (EP’), and $(\rho(m),v)$ satisfies (EP) (Part (iii)). Parts (iv) and (v) come from Remark 5.16, Remark 5.17, Corollary 5.24 and Corollary 5.26 (Part (iv) and Part (v)). ∎
@@
The proof of Theorem 2.13 is essentially the same as the proof of Theorem 2.12. Two things we need to note are (1) if we don’t know $\gamma>2$, we may not have the fact that there is a “uniform bound on the radius of $\rho(m)$’s support for all small mm” from Corollary 5.24. However, since the radii of $\Omega_{1-m}$ and $\Omega_{m}$ is $\frac{\eta}{4}\sim m^{-2}$, recall $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$, due to Corollary 5.24, we can still have the fact that $\rho(m)$ is located in the interior of $\Omega_{m}\cup\Omega_{1-m}$ if ${\frac{\gamma-2}{3\gamma-4}}>-2$, that is $\gamma>\frac{10}{7}$; (2) in order to obtain the energy convergence in Proposition 5.8, we hope the term $Cm^{\frac{4\gamma-6}{3\gamma-4}}$ in (39) satisfies $\frac{4\gamma-6}{3\gamma-4}>0$, which means $\gamma>\frac{3}{2}$. Therefore, when $\gamma>\max\{\frac{3}{2},\frac{10}{7}\}=\frac{3}{2}$, we have Theorem 2.13 holds true. ∎
@@ 7 Upper Bounds for Distances Between Connected Components of Minimizers
It should be noted that, although we prove the existence of Wasserstein $L^{\infty}$ local minimizer on $R(\mathbb{R}^{3}$), which is also a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ on $W_{m}$ (Theorem 2.12 and Theorem 2.13), in particular, spt $\rho(m)$ has at least two connected components, it is still unclear whether it has exactly two connected components. Neither we in this paper nor McCann [46, Theorem 6.1 and Corollary 6.2] elaborated on or provided proof for the result. If spt $\rho_{1-m}$ or spt $\rho_{m}$ have multiple connected components, the model may correspond to a system with multiple stars and planets, such as the relationship between the Solar System and the asteroid belt. Moreover, one could interpret it as describing a large and a small galaxy; therefore the existence results retain physical relevance and may inspire further study of the Vlasov–Poisson system via Rein’s reduction method (see e.g. [31, 49, 50]).
@@
Then we know $d=\inf\limits_{y\in V}y_{1}-\sup\limits_{x\in U}x_{1}$. We denote the mass in $U$ is $m_{1}\coloneq\int_{U}\rho_{m}$, then the mass in $V$ is $m_{2}\coloneq\int_{V}\rho_{m}=m-m_{1}$. Notice $\rho_{m}(x)=\rho_{U}(x)+\rho_{V}(x)$, where $\rho_{U}(x)=\rho_{m}(x)\cdot\mathbf{1}_{U}(x)$, $\rho_{V}(x)=\rho_{m}(x)\cdot\mathbf{1}_{V}(x)$, $\mathbf{1}_{E}$ is the indicator function of set $E$. We want to move $\rho_{U}(x)$ and $\rho_{V}(x)$ in the following way: given $h>0$, let $h_{1}>0$ and $h_{2}>0$ satisfy $h_{1}+h_{2}=h$, $h_{1}\cdot m_{1}=h_{2}\cdot m_{2}$. With an abuse of notation we denote $h_{1}=h_{1}\cdot\hat{e_{1}}=(h_{1},0,0)^{T}$, similarly $h_{2}=(h_{2},0,0)^{T}$, $h=(h,0,0)^{T}$. Let $\rho_{m}^{[h]}=\rho_{U}(x-h_{1})+\rho_{V}(x+h_{2})$. Easy to see when $h$ is small enough, the connected components of spt $\rho_{m}^{[h]}$ are $U_{h}=U+h_{1}$, and $V_{h}=V-h_{2}$, and we still have $\sup\limits_{x\in U_{h}}x_{1}<b<\inf\limits_{y\in V_{h}}y_{1}$. We can write $\rho_{m}^{[h]}=\rho_{U}^{[h]}+\rho_{V}^{[h]}$ where $\rho_{U}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{U_{h}}$ and $\rho_{V}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{V_{h}}$. We notice the centers of mass of $\rho_{m}$ and $\rho_{m}^{[h]}$ are the same. Let $\rho^{[h]}(m)=\rho_{m}^{[h]}+\rho_{1-m}$, then we will show if $d$ is too large, then $E_{J}(\rho^{[h]}(m))<E_{J}(\rho(m))$.
@@
$\begin{split}I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)=\frac{m_{1}m_{2}}{m}\left({-2h\left({{\bar{x}}_{1}\left(\rho_{V}\right)-{\bar{x}}_{1}\left(\rho_{U}\right)}\right)+h^{2}}\right)=\frac{1}{m}{\int{-2h\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)d^{3}xd^{3}y}}+\frac{m_{1}m_{2}}{m}h^{2}\end{split}$
@@
$\begin{split}&\quad T_{J}\left({\rho(m)}\right)-T_{J}\left({\rho^{[h]}(m)}\right)\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)}-\frac{J^{2}}{2I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}\left({I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)}\right)}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}\cdot{\left({{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{-2h\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)d^{3}xd^{3}y}}+m_{1}m_{2}h^{2}}\right)}\end{split}$ (44)
@@
$\begin{split}G\left({\rho_{m},\rho_{m}}\right)-G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)&=G\left({\rho_{U}+\rho_{V},\rho_{U}+\rho_{V}}\right)-G\left({\rho_{U}^{[h]}+\rho_{V}^{[h]},\rho_{U}^{[h]}+\rho_{V}^{[h]}}\right)\\ &=2G\left({\rho_{U},\rho_{V}}\right)-2G\left({\rho_{U}^{[h]},\rho_{V}^{[h]}}\right)\\ &=2{\int{\frac{\rho_{U}(x)\rho_{V}(y)-\rho_{U}\left({x-h_{1}}\right)\rho_{V}\left({y+h_{2}}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &=2{\int{\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y}\right|}-\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y+h}\right|}d^{3}xd^{3}y}}\\ &=2{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)}\right)d^{3}xd^{3}y}}\\ &=2{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}d^{3}xd^{3}y}}+2m_{1}m_{2}O\left(h^{2}\right)\\ &\leq-\frac{2}{R^{3}(m)}{\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}+2m_{1}m_{2}O\left(h^{2}\right)\end{split}$ (46)
@@
$\begin{split}&G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\\ &={\int{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)+\rho_{V}(y)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}-{\int{\frac{\rho_{1-m}(x)\left({\rho_{U}\left({y-h_{1}}\right)+\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &={\int{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)-\rho_{U}\left({y-h_{1}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}+{\int{\frac{\rho_{1-m}(x)\left({\rho_{V}(y)-\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &={\int{\rho_{1-m}(x)\rho_{U}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y-h_{1}}\right|}}\right)d^{3}xd^{3}y}}+{\int{\rho_{1-m}(x)\rho_{V}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h_{2}}\right|}}\right)d^{3}xd^{3}y}}\\ &\leq{\int{\rho_{1-m}(x)\rho_{U}(y)\left(\frac{h_{1}}{\left|{x-y}\right|\left|{x-y-h_{1}}\right|}\right)d^{3}xd^{3}y}}+{\int{\rho_{1-m}(x)\rho_{V}(y)\left(\frac{h_{2}}{\left|{x-y}\right|\left|{x-y+h_{2}}\right|}\right)d^{3}xd^{3}y}}\end{split}$
By the construction of $\Omega_{1-m}$ and $\Omega_{m}$ (23) we know $\rho_{1-m}(x)>0$, $\rho_{m}(y)>0$ implies $\frac{J^{2}}{2m^{2}\left({1-m}\right)^{2}}\leq\left|{x-y}\right|\leq\frac{3J^{2}}{2m^{2}\left({1-m}\right)^{2}}$, and when $h$ is small enough, $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{1}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$ and $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{2}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$. We also notice that by construction ${\int{\rho_{U}h_{1}}}={\int{\rho_{V}h_{2}}}=\frac{m_{1}m_{2}h}{m_{1}+m_{2}}=\frac{m_{1}m_{2}h}{m}$, then
@@
$\begin{split}&E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}(m)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}}\right)\cdot{\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}(m)}-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$ (48)
Therefore, when $d>\frac{32m^{3}\left({1-m}\right)^{5}R^{3}(m)}{J^{4}}=\frac{32\left({1-m}\right)^{5}R^{3}(J)}{J^{4}}m^{\frac{12\gamma-18}{3\gamma-4}}$, we know $E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$ for $h$ small enough. That is, if $\rho(m)$ is a constrained minimizer, then $d\leq Cm^{\frac{12\gamma-18}{3\gamma-4}}$ for some $C$ and all small $m$. More precisely, we have the following proposition:
@@
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13. There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{m}$ has 2 convex connected components, then the distance between these 2 components is less then $\eta$, where $\eta=Cm^{\frac{12\gamma-18}{3\gamma-4}}$.

### v2

@@ 6.2 Existence theorem
After making some modifications to McCann’s proof in [24, Section 6], we proceed by translating $\rho$ such that it enjoys a plane of symmetry $z=0$, and then use the results above to show $\rho$ is supported away from the boundary of $\Omega_{1-m}\cup\Omega_{m}$, then we finalize the proof of existence.
@@
Take $m$ small enough so that minimizers exist. First, we claim that any minimizer $\rho(m)=\rho_{m}+\rho_{1-m}$ for $E_{J}(\rho)$ on $W_{m}$ may be translated so that both $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=0$. In fact, since the $\Omega_{m}$ and $\Omega_{1-m}$ are convex and symmetric about $z=0$, it is enough to know that $\rho(m)$ enjoys a plane of symmetry $z=c$. This follows from a strong rearrangement inequality in Lieb [21, Lemma 3] and Fubini’s Theorem: by the definition of $U(\rho)$ and $I(\rho)$, they will not be changed after the symmetric decreasing rearrangement of $\rho$ along lines parallel to the z-axis; however, since $\left(r^{2}+z^{2}\right)^{-\frac{1}{2}}$ is strictly decreasing as a function of $|z|$, the rearrangement increases $G(\rho,\rho)$ unless $\rho$ is already symmetric decreasing about a plane $z=c$. Note $\Omega_{1-m}$ and $\Omega_{m}$ are convex, hence $\rho(m)$’s rearrangement is in $W_{m}$. Since $\rho(m)$ minimizes $E_{J}(\rho)$, $G(\rho,\rho)$ cannot be increased after rearrangement, which means $\rho(m)$ enjoys a plane of symmetry $z=c$.
@@
Let $\delta$ as in Lemma 6.3 and $\rho(m)$ be the constrained energy minimizer in Theorem 4.14, then Lemma 6.3 shows that spt $\rho({m})$ is compact in the interior of $\Omega_{m}\cup\Omega_{1-m}$, therefore separated from the boundary by a positive distance $\zeta$. Lemma 2.9 (ii) shows that if $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho(m),\kappa)<\eta$, then $\kappa$ in $W_{m}$. Thus $E_{J}(\rho(m))\leq E_{J}(\kappa)$, and $\rho(m)$ is a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$ (Existence of $W^{\infty}$ local energy minimizer). By the definition of $W_{m}$ (22), we know all the functions in $W_{m}$ have compact support. Thus after translation we can set $\rho(m)\in{R}_{0}\left(\mathbb{R}^{3}\right)$. Since $\rho(m)$ is a $W^{\infty}$ local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$, in particular on ${R}_{0}\left(\mathbb{R}^{3}\right)$, Theorem 3.1 provides a local minimizer $(\rho(m),v)$ of $E(\rho,v)$ subject to the constraint on $J_{z}=J$, with $v(x)=$ $\frac{J}{I(\rho(m))}\hat{e}_{z}\times x$. [24, Section 2] and [10, Section 2] show that $(\rho(m),v)$ minimizes $E(\rho,v)$ subject to the constraint on the vector angular momentum as well (Part (i)). Part (ii) comes from Lemma 6.3 (Part (ii)). By Theorem 3.1 we know $\rho(m)$ is continuous and satisfies (EP’), and $(\rho(m),v)$ satisfies (EP) (Part (iii)). Parts (iv) and (v) come from Remark 5.16, Remark 5.17, Corollary 5.23 and Corollary 5.25 (Part (iv) and Part (v)). ∎
@@
The proof of Theorem 2.13 is essentially the same as the proof of Theorem 2.12. Two things we need to note are (1) if we do not know $\gamma>2$, we may not have the fact that there is a “uniform bound on the radius of $\rho(m)$’s support for all small mm” from Corollary 5.23. However, since the radii of $\Omega_{1-m}$ and $\Omega_{m}$ is $\frac{\eta}{4}\sim m^{-2}$, recall $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$, due to Corollary 5.23, we can still have the fact that $\rho(m)$ is located in the interior of $\Omega_{m}\cup\Omega_{1-m}$ if ${\frac{\gamma-2}{3\gamma-4}}>-2$, that is $\gamma>\frac{10}{7}$; (2) in order to obtain the energy convergence in Proposition 5.8, we hope the term $Cm^{\frac{4\gamma-6}{3\gamma-4}}$ in (40) satisfies $\frac{4\gamma-6}{3\gamma-4}>0$, which means $\gamma>\frac{3}{2}$. Therefore, when $\gamma>\max\{\frac{3}{2},\frac{10}{7}\}=\frac{3}{2}$, Theorem 2.13 holds. ∎
@@ 7 Upper Bounds for Distances Between Connected Components of Minimizers
It should be noted that, although we prove the existence of Wasserstein $L^{\infty}$ local minimizer on $R(\mathbb{R}^{3}$), which is also a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ on $W_{m}$ (Theorem 2.12 and Theorem 2.13), in particular, spt $\rho(m)$ has at least two connected components, it is still unclear whether it has exactly two connected components. Neither we in this paper nor McCann [24, Theorem 6.1 and Corollary 6.2] elaborated on or provided proof for the result. If spt $\rho_{1-m}$ or spt $\rho_{m}$ have multiple connected components, the model may correspond to a system with multiple stars and planets, such as the relationship between the Solar System and the asteroid belt. Moreover, one could interpret it as describing a large and a small galaxy; therefore the existence results retain physical relevance and may inspire further study of the Vlasov–Poisson system via Rein’s reduction method (see e.g. [17, 26, 27]).
@@
Then we know $d=\inf\limits_{y\in V}y_{1}-\sup\limits_{x\in U}x_{1}$. We denote the mass in $U$ by $m_{1}\coloneq\int_{U}\rho_{m}\,dx$, then the mass in $V$ is $m_{2}\coloneq\int_{V}\rho_{m}\,dx=m-m_{1}$. Notice $\rho_{m}(x)=\rho_{U}(x)+\rho_{V}(x)$, where $\rho_{U}(x)=\rho_{m}(x)\cdot\mathbf{1}_{U}(x)$, $\rho_{V}(x)=\rho_{m}(x)\cdot\mathbf{1}_{V}(x)$, $\mathbf{1}_{E}$ is the indicator function of set $E$. We want to move $\rho_{U}(x)$ and $\rho_{V}(x)$ in the following way: given $h>0$, let $h_{1}>0$ and $h_{2}>0$ satisfy $h_{1}+h_{2}=h$ and $h_{1}\cdot m_{1}=h_{2}\cdot m_{2}$. With an abuse of notation we denote $h_{1}=h_{1}\cdot\hat{e_{1}}=(h_{1},0,0)^{T}$, similarly $h_{2}=(h_{2},0,0)^{T}$, $h=(h,0,0)^{T}$. Then we define $\rho_{m}^{[h]}=\rho_{U}(x-h_{1})+\rho_{V}(x+h_{2})$.
Easy to see when $h$ is small enough, the connected components of spt $\rho_{m}^{[h]}$ are $U_{h}=U+h_{1}$ and $V_{h}=V-h_{2}$, and we still have $\sup\limits_{x\in U_{h}}x_{1}<b<\inf\limits_{y\in V_{h}}y_{1}$. We can write $\rho_{m}^{[h]}=\rho_{U}^{[h]}+\rho_{V}^{[h]}$ where $\rho_{U}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{U_{h}}$ and $\rho_{V}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{V_{h}}$. We notice the centers of mass of $\rho_{m}$ and $\rho_{m}^{[h]}$ are the same.
Let $\rho^{[h]}(m)=\rho_{m}^{[h]}+\rho_{1-m}$, then we will show if $d$ is too large, then $E_{J}(\rho^{[h]}(m))<E_{J}(\rho(m))$.
@@
$\begin{split}I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)&=\frac{m_{1}m_{2}}{m}\left({-2h\left({{\bar{x}}_{1}\left(\rho_{V}\right)-{\bar{x}}_{1}\left(\rho_{U}\right)}\right)+h^{2}}\right)\\ &=\frac{1}{m}{(-2h)\iint{\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)\,dx\,dy}}+\frac{m_{1}m_{2}}{m}h^{2}\end{split}$
@@
$\begin{split}&\quad T_{J}\left({\rho(m)}\right)-T_{J}\left({\rho^{[h]}(m)}\right)\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)}-\frac{J^{2}}{2I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}\left({I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)}\right)}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}\cdot{\left({{(-2h)\iint_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)\,dx\,dy}}+m_{1}m_{2}h^{2}}\right)}\end{split}$ (45)
@@
$\begin{split}G\left({\rho_{m},\rho_{m}}\right)-G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)&=G\left({\rho_{U}+\rho_{V},\rho_{U}+\rho_{V}}\right)-G\left({\rho_{U}^{[h]}+\rho_{V}^{[h]},\rho_{U}^{[h]}+\rho_{V}^{[h]}}\right)\\ &=2G\left({\rho_{U},\rho_{V}}\right)-2G\left({\rho_{U}^{[h]},\rho_{V}^{[h]}}\right)\\ &=2{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{U}(x)\rho_{V}(y)-\rho_{U}\left({x-h_{1}}\right)\rho_{V}\left({y+h_{2}}\right)}{\left|{x-y}\right|}\,dx\,dy}}\\ &=2{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y}\right|}-\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y+h}\right|}\,dx\,dy}}\\ &=2{\iint_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)}\right)\,dx\,dy}}\\ &=2h{\iint_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\frac{\left(-\left({y_{1}-x_{1}}\right)\right)}{\left|{x-y}\right|^{3}}\,dx\,dy}}+2m_{1}m_{2}O\left(h^{2}\right)\\ &\leq-\frac{2h}{R^{3}(m)}{\iint_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)\,dx\,dy}}+2m_{1}m_{2}O\left(h^{2}\right)\end{split}$ (47)
@@
$\begin{split}&G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\\ &={\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)+\rho_{V}(y)}\right)}{\left|{x-y}\right|}\,dx\,dy}}-{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{U}\left({y-h_{1}}\right)+\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}\,dx\,dy}}\\ &={\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)-\rho_{U}\left({y-h_{1}}\right)}\right)}{\left|{x-y}\right|}\,dx\,dy}}+{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{V}(y)-\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}\,dx\,dy}}\\ &={\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{U}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y-h_{1}}\right|}}\right)\,dx\,dy}}+{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{V}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h_{2}}\right|}}\right)\,dx\,dy}}\\ &\leq{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{U}(y)\left(\frac{h_{1}}{\left|{x-y}\right|\left|{x-y-h_{1}}\right|}\right)\,dx\,dy}}+{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{V}(y)\left(\frac{h_{2}}{\left|{x-y}\right|\left|{x-y+h_{2}}\right|}\right)\,dx\,dy}}\end{split}$
By the construction of $\Omega_{1-m}$ and $\Omega_{m}$ (24) we know $\rho_{1-m}(x)>0$, $\rho_{m}(y)>0$ implies $\frac{J^{2}}{2m^{2}\left({1-m}\right)^{2}}\leq\left|{x-y}\right|\leq\frac{3J^{2}}{2m^{2}\left({1-m}\right)^{2}}$, and when $h$ is small enough, $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{1}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$ and $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{2}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$. We also notice that by construction
${\int_{\mathbb{R}^{3}}{\rho_{U}h_{1}}\,dx}={\int_{\mathbb{R}^{3}}{\rho_{V}h_{2}}\,dx}=\frac{m_{1}m_{2}h}{m_{1}+m_{2}}=\frac{m_{1}m_{2}h}{m}$
Then
@@
$\begin{split}&E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}(m)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}}\right)\cdot h\cdot{\iint_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)\,dx\,dy}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}(m)}-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$ (49)
Therefore, when $d>\frac{32m^{3}\left({1-m}\right)^{5}R^{3}(m)}{J^{4}}=\frac{32\left({1-m}\right)^{5}R^{3}(J)}{J^{4}}m^{\frac{12\gamma-18}{3\gamma-4}}$, we know for $h$ small enough,
$E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$
That is, if $\rho(m)$ is a constrained minimizer, then $d\leq Cm^{\frac{12\gamma-18}{3\gamma-4}}$ for some $C$ and all small $m$. More precisely, we have the following proposition:
@@
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13. There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{m}$ has 2 convex connected components, then the distance between these 2 components is less than $\eta$, where $\eta=Cm^{\frac{12\gamma-18}{3\gamma-4}}$.